Find the first two terms and the 50th term of each infinite sequence. An=2n-3
Answer by jim_thompson5910(35256)

First term:

An=2n-3

A1=2(1)-3

A1=2-3

A1=-1

-------------------------------------------------------

Second term:

An=2n-3

A2=2(2)-3

A2=4-3

A2=1

-------------------------------------------------------
50th term:

An=2n-3

A50=2(50)-3

A50=100-3

A50=97

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------